Determine p- and o- from the given parameters of the population and sample size. = 71, o = 16, n = 64
Answer #1

X ~ N ( µ = 71 , σ = 16 )

Sampling distribution of sample mean

Mean of sampling distribution µ = µ = 71


Standard deviation of sample mean σ = σ / √ (n) = 16/√64 = 2